OpenStudy (anonymous):

I am such a ditz!!!! ln (4x-7)= ln (x+11). Please I need help!!!!

OpenStudy (anonymous):

since natural log is being taken of both sides you can just equate whats inside 4x-7=x+11

OpenStudy (anonymous):

DIFFERENTIATE on but sides, it will also lead you to a perfect answer

terenzreignz (terenzreignz):

Because \[\huge \ln a = \ln b \rightarrow e^{\ln a} = e^{\ln b}\rightarrow a=b\]

terenzreignz (terenzreignz):

@soty2013 That may not be feasible... You'll get \[\frac{4}{4x-7}=\frac{1}{x+11}\] I don't think this gives you the answer you want...

Directrix (directrix):

What value of x makes this true: 4x-7=x+11? @dawnd4fun